# Nonharmonic multivariate Fourier transforms and matrices: condition numbers and hyperplane geometry

### v1

@@ Abstract
Consider an operator that takes the Fourier transform of a discrete measure supported in $\mathcal{X}\subseteq[-\frac{1}{2},\frac{1}{2})^{d}$ and restricts it to a compact $\Omega\subseteq\mathbb{R}^{d}$. We provide lower bounds for its smallest singular value when $\Omega$ is either a ball or cube of radius $m$, and under different types of geometric assumptions on $\mathcal{X}$. We first show that if distances between points in $\mathcal{X}$ are lower bounded by a $\delta$ that is allowed to be arbitrarily small, then the smallest singular value is at least $Cm^{d/2}(m\delta)^{\lambda-1}$, where $\lambda$ is the maximum number of elements in $\mathcal{X}$ contained within any ball or cube of an explicitly given radius. This estimate communicates a localization effect of the Fourier transform. While it is sharp, the smallest singular value behaves better than expected for many $\mathcal{X}$, including when we dilate a generic set by parameter $\delta$. We next show that if there is a $\eta$ such that, for each $x\in\mathcal{X}$, the set $\mathcal{X}\setminus\{x\}$ locally consists of at most $r$ hyperplanes whose distances to $x$ are at least $\eta$, then the smallest singular value is at least $Cm^{d/2}(m\eta)^{r}$. For dilations of a generic set by $\delta$, the lower bound becomes $Cm^{d/2}(m\delta)^{\lceil(\lambda-1)/d\rceil}$. The appearance of a $1/d$ factor in the exponent indicates that compared to worst case scenarios, the condition number of nonharmonic Fourier transforms is better than expected for typical sets and improve with higher dimensionality.
@@ 1.1 Motivation
In the univariate case, it is well understood that the smallest singular value greatly depends on $m$, $s$, and configuration of $\mathcal{X}$, see [23, 9, 39, 2, 35, 6, 30, 7, 34]. On the other hand, the maximum singular value is not as interesting to study. Here and throughout this paper, we let $|\Omega|$ be the Lebesgue measure of $\Omega$, while we $|\Omega|_{*}=|\Omega\cap\mathbb{Z}^{d}|$ be the number of multi-integers contained in $\Omega$. It is not hard to show that
@@ 1.2 Some geometric concepts
We say $\mathcal{X}$ consists of clumps with parameters $(p,\tau,\lambda)$ if the following hold. There exist $r\in\mathbb{N}_{+}$ and disjoint subsets $\mathcal{C}_{1},\dots,\mathcal{C}_{r}$ of $\mathcal{X}$ called clumps such that $\mathcal{X}=\mathcal{C}_{1}\cup\cdots\cup\mathcal{C}_{r}$, the $\ell^{p}$ distance between any two elements in the same clump is bounded above by $\tau$, and $|\mathcal{C}_{k}|=\lambda$. Further, if $r>1$, assume that the $\ell^{p}$ distance between different $\mathcal{C}_{j}$ and $\mathcal{C}_{k}$ exceeds $\tau$.
Aside from clumps, a more complicated order geometric description can be given in terms of hyperplane decompositions. We say a set $\mathcal{X}$ consists of $r$ hyperplanes relative to a $x\in\mathcal{X}$ (called the reference point) if there exist hyperplanes $\mathcal{H}_{1},\dots,\mathcal{H}_{r}\subseteq\mathbb{R}^{d}$ that do not intersect $x$ and
@@ 1.3 Summary of main results and organization
There are many applications where a strong separation condition is violated. In Section 3, we focus on the super-resolution regime, which allows for $\Delta_{p^{\prime}}(\mathcal{X})$ to be arbitrarily small. As the separation goes to zero, the smallest singular values of both Fourier operators tend to zero, and our primary goal is to derive non-asymptotic and sharp lower bounds. Theorems 3.9 and 3.10 are multiscale estimates which show that the distances between elements in $\mathcal{X}$ at small scales predominately influences these operators’ smallest singular values. We specialize these theorems to clumps in order to give more interpretable results. Corollaries 3.11 and 3.12 show that if $\mathcal{X}$ consists of clumps with parameters $(p,\tau,\lambda)$ where $\tau\gtrsim\frac{\lambda}{m}$ and $\Delta_{p^{\prime}}(\mathcal{X})$ is sufficiently small, then
@@ 2 Well-separated case
Let $\Omega\subseteq\mathbb{R}^{d}$ be compact, $p\in[1,\infty]$, and $\delta\in(0,\frac{1}{2})$. Suppose there exist $\varphi$ and $\psi$ that are supported in $\Omega_{\delta}^{p}$ and $\widehat{\psi}(\xi)\leq\mathbbm{1}_{\Omega}(\xi)\leq\widehat{\varphi}(\xi)$ for all $\xi\in\mathbb{R}^{d}$. For any set $\mathcal{X}\subseteq[-\frac{1}{2},\frac{1}{2})^{d}$ such that $\Delta_{p}(\mathcal{X})>\delta$, we have
@@
For any $\delta>0$ and integer $d\geq 2$, the following hold. There exist $\psi$ and $\varphi$ such that $\widehat{\psi}(\xi)\leq\mathbbm{1}_{Q_{1}}(\xi)\leq\widehat{\varphi}(\xi)$ for all $\xi\in\mathbb{R}^{d}$, they are supported in $Q_{\delta}$, and
$\int_{\mathbb{R}^{d}}(\widehat{\varphi}-\mathbbm{1}_{Q_{1}})=\int_{\mathbb{R}^{d}}(\mathbbm{1}_{Q_{1}}-\widehat{\psi})$ $=\Big{(}\big{(}1+\tfrac{1}{2\delta}\big{)}^{d}-1\Big{)}\,|Q_{1}|.$
@@
$\psi(0)$ $=|Q_{1}|-\int_{\mathbb{R}^{d}}(\mathbbm{1}_{Q_{1}}-\widehat{\psi})=\Big{(}2-\big{(}1+\tfrac{1}{2\delta}\big{)}^{d}\Big{)}\,|Q_{1}|>(2-e^{d/(2\delta)})|Q_{1}|\geq 0,$ (2.1)
@@
Notice that the $\ell^{\infty}$ separation condition for $\mathcal{X}$ is inversely proportional to $m$, which is expected from scaling arguments. Additionally, the constant $\beta d$ grows linearly in the dimension $d$, but it not clear whether this growth on $d$ is necessary.
Let us compare Theorem 2.3 with Kunis-Nagel-Strotmann [31, Theorem 4.2]. They parameterized the frequency set by $\{0,\dots,N-1\}^{d}$, while we use $Q_{m}=\{-m,\dots,m\}^{d}$. The singular values of $\Phi_{\Omega,\mathcal{X}}$ are invariant under multi-integer shifts of $\Omega$, so our settings are equivalent whenever we enforce the relationship $N=2m+1$. They proved that $\Delta_{2}(\mathcal{X})>\frac{4d}{2m+1}$ implies $\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})>0.9\,\sqrt{N^{d}}=0.9\sqrt{|Q_{m}|_{*}}$. To get the same conclusion, we pick $\beta=(2\log(2-0.81))^{-1}\approx 0.7671$ in Theorem 2.3 to see that $\Delta_{2}(\mathcal{X})\geq\frac{\beta d}{m}$ implies $\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})\geq 0.9\,\sqrt{|Q_{m}|_{*}}$. Both results require separation conditions that grow linearly in $d$, but our universal constant is smaller due to Barton.
@@
$c(\alpha):=\frac{1}{|\mathbb{S}^{d-1}|}\Big{(}\frac{2}{\alpha}\Big{)}^{d}\frac{\gamma(\alpha)}{1+\gamma(\alpha)/d}\quad\text{and}\quad\gamma(\alpha):=\frac{-\pi\alpha J_{d/2-1}(\pi\alpha)}{J_{d/2}(\pi\alpha)}.$ (2.3)
@@
Let $d\geq 2$ be an integer and $\alpha\in(\frac{j_{d/2-1,1}}{\pi},\frac{j_{d/2,1}}{\pi})$. Then there exists $\psi\colon\mathbb{R}^{d}\to\mathbb{R}$ supported in $B_{\alpha}$ such that $\widehat{\psi}(\xi)\leq\mathbbm{1}_{B_{1}}(\xi)$ for all $\xi\in\mathbb{R}^{d}$ and $\psi(0)=c(\alpha).$ If $\alpha\leq\frac{j_{d/2-1,1}}{\pi}$ there is no $\psi$ satisfying these properties with $\psi(0)>0$.
@@
$c(\alpha):=\frac{1}{|\mathbb{S}^{d-1}|}\Big{(}\frac{2\pi}{j_{d/2,1}}\Big{)}^{d}\quad\text{if}\quad\alpha\geq\frac{j_{d/2,1}}{\pi}.$ (2.4)
@@
There are related but not directly comparable results. For $d=2$, Chen-Moitra [14] studied how small $\Delta_{2}(\mathcal{X})$ can be so that $\sigma_{\min}(\mathcal{F}_{B_{1/2},\mathcal{X}})$ is not exponentially small in $s$. By selecting $\mathcal{X}$ to be $s$ points on a hexagonal lattice of $\mathbb{R}^{2}$ and letting $s$ grow, they showed that $\Delta_{2}(\mathcal{X})\geq\sqrt{4/3}\approx 1.1547$ is required for stable recovery of $\mu$, in some appropriate sense defined in their paper. This construction can be adapted to derive an upper bound (as opposed to lower) for $\sigma_{\min}(\mathcal{F}_{B_{1/2},\mathcal{X}})$. Their conclusion that a minimum separation of $\sqrt{4/3}$ is necessary results in allowing $s$ to arbitrarily large. We will see in the next section that provided $s$ is fixed, the minimum separation can be arbitrarily small.
@@ 3 Super-resolution regime and arbitrary nodes
In the super-resolution regime, a primary method that we will use to lower bound $\sigma_{\min}(\mathcal{F}_{\Omega,\mathcal{X}})$ and $\sigma_{\min}(\Phi_{\Omega,\mathcal{X}})$ is through a “dual” relationship with minimum norm interpolation. This duality was introduced in [35, Proposition 2.12] for dealing with $\Phi_{\Omega,\mathcal{X}}$ in the one dimensional case. It readily generalizes to higher dimensions and to $\mathcal{F}_{\Omega,\mathcal{X}}$ with minor modifications. This dual relationship was further combined with density ideas and sparsity decompositions in [34], which significantly improved existing estimates for one-dimensional Fourier matrices.
We first collect many these ideas before moving onto the main theorems. While we provide some exposition, it is rather brief; for those interested in the intuition and concepts behind these techniques, we refer the reader to [34], which explains them in greater detail and in the simpler one-dimensional case. Recall that throughout, we assume $\Omega\subseteq\mathbb{R}^{d}$ is compact. We let $\mathcal{W}(\Omega)$ denote the space of bandlimited functions or the Paley-Wiener space; that is, a complex-valued $f$ belongs to $\mathcal{W}(\Omega)$ if and only if there is a $F\in L^{2}(\Omega)$ such that
@@
This is a straightforward adaptation of [35, Proposition 2.12], but we include its proof in Section 6.2 for completeness. Typically, we will use this proposition for the case when $\varepsilon=0$ and $v$ is a right singular vector of $\mathcal{F}_{\Omega,\mathcal{X}}$ (or $\Phi_{\Omega,\mathcal{X}}$) corresponding to its smallest singular value. It is prudent to mention that this proposition does not make any claims about the relationship between singular values $\mathcal{F}_{\Omega,\mathcal{X}}$ and $\Phi_{\Omega,\mathcal{X}}$ because they in general have different right singular vectors, so this statement would be applied to different $v$’s.
@@
The next proposition acts as a converse to Proposition 3.1, and shows that any lower bound on the smallest singular value provides the existence of polynomials with prescribed interpolation properties. A proof of the following lemma for $d=1$ can be found in [34, Proposition 5.3] and readily extends to higher dimensions.
@@
This proposition can be readily deduced from a greedy construction of the $\mathcal{X}_{1},\dots,\mathcal{X}_{\nu}$, see the proof of [34, Proposition 6.1] for $d=1$. The next proposition illustrates a nontrivial application of the polynomial method and is proved in Section 6.4.
@@
$\sigma_{\min}(\Phi_{\Omega^{p}_{C/\tau},\mathcal{W}})\geq c\sqrt{|\Omega^{p}_{C/\tau}|_{*}}.$ (3.2)
@@
Provided that the assumptions hold, this localization comes at a price of course. Comparing both sides of the inequality in (3.3), the number of Fourier samples are reduced from $|\Omega^{p}_{m}|_{*}$ to $|\Omega^{p}_{m/2}|_{*}$, so resolution worsens by a factor of 2. We also pick up an additional $c^{\nu_{p}(\tau,\mathcal{X})}$ term, which can be small if either $c$ is close to zero and/or $\nu_{p}(\tau,\mathcal{X})$ is large. For many sets of interest, the $\tau$ local sparsity of $\mathcal{X}$ is small. When we apply this theorem, the constant $c$ will come from Theorems 2.5 and 2.3, so it is important to obtain good constants for the well separated case to temper the effects of localization.
With the localization lemma at hand, let us explain at a high level how to deal with the local set $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$. We will again employ the duality principle, which transforms the problem of constructing Lagrange polynomials $\{f_{k}\}_{k=1}^{s}$ for $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$. Shifting by $x_{k}$ and noting that both $\mathcal{P}(\Omega)$ and $\mathcal{W}(\Omega)$ are invariant under shifts, this reduces the problem down to the canonical case whereby $0\in\mathcal{U}$ and all other elements in $\mathcal{U}$ are sufficiently close to zero. Importantly, we place no lower bounds on $\Delta_{p}(\mathcal{U})$. We seek a $f$ such that $f(0)=1$, $f$ vanishes on $\mathcal{U}\setminus\{0\}$, and its $L^{2}$ norm is suitably controlled.
@@
Second, we need to pick $q(u)$ carefully so that $|q(u)\cdot u|$, in order to ensure that $\|\widetilde{f}\|_{L^{\infty}(\mathbb{T}^{d})}$ is reasonably controlled, yet $|q(u)|_{p}$ itself is not too large otherwise the obtained interpolant might not be in $\mathcal{P}(\Omega)$. See Figure 2 for a cartoon of how replacing $a$ with $q$ effectively changes the direction of $f$.
@@
$\sigma_{\min}(\mathcal{F}_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{s}}\Big{(}\frac{|B_{1}|}{c(\alpha)}\Big{)}^{\nu/2}\Big{(}\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\Big{)}^{\nu/2}\,\sqrt{|B_{m/(2\nu)}|}\,\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{2}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{2}\Bigg{\}},$
$\sigma_{\min}(\Phi_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{s}}\Big{(}\frac{|B_{1}|}{c(\alpha)}\Big{)}^{\nu/2}\Big{(}\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\Big{)}^{\nu/2}\sqrt{|B_{m/(2\nu)}|_{*}}\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{2}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{2}\Bigg{\}}.$
@@
$\sigma_{\min}(\mathcal{F}_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\Big{(}1-\frac{1}{2}e^{1/(2\beta)}\Big{)}^{\nu/2}\,\sqrt{|Q_{m/(2\nu)}|}\,\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{1}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{1}\Bigg{\}},$
$\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\Big{(}1-\frac{1}{2}e^{1/(2\beta)}\Big{)}^{\nu/2}\sqrt{|Q_{m/(2\nu)}|_{*}}\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{1}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{1}\Bigg{\}}.$

### v2

@@ Abstract
Consider an operator that takes the Fourier transform of a discrete measure supported in $\mathcal{X}\subseteq[-\frac{1}{2},\frac{1}{2})^{d}$ and restricts it to a compact $\Omega\subseteq\mathbb{R}^{d}$. We provide lower bounds for its smallest singular value when $\Omega$ is either a closed ball of radius $m$ or closed cube of side length $2m$, and under different types of geometric assumptions on $\mathcal{X}$. We first show that if distances between points in $\mathcal{X}$ are lower bounded by a $\delta$ that is allowed to be arbitrarily small, then the smallest singular value is at least $Cm^{d/2}(m\delta)^{\lambda-1}$, where $\lambda$ is the maximum number of elements in $\mathcal{X}$ contained within any ball or cube of an explicitly given radius. This estimate communicates a localization effect of the Fourier transform. While it is sharp, the smallest singular value behaves better than expected for many $\mathcal{X}$, including when we dilate a generic set by parameter $\delta$. We next show that if there is a $\eta$ such that, for each $x\in\mathcal{X}$, the set $\mathcal{X}\setminus\{x\}$ locally consists of at most $r$ hyperplanes whose distances to $x$ are at least $\eta$, then the smallest singular value is at least $Cm^{d/2}(m\eta)^{r}$. For dilations of a generic set by $\delta$, the lower bound becomes $Cm^{d/2}(m\delta)^{\lceil(\lambda-1)/d\rceil}$. The appearance of a $1/d$ factor in the exponent indicates that compared to worst case scenarios, the condition number of nonharmonic Fourier transforms is better than expected for typical sets and improve with higher dimensionality.
@@ 1.1 Motivation
In the univariate case, it is well understood that the smallest singular value greatly depends on $m$, $s$, and configuration of $\mathcal{X}$, see [23, 9, 41, 2, 36, 6, 31, 7, 35]. On the other hand, the maximum singular value is not as interesting to study. Here and throughout this paper, we let $|\Omega|$ be the Lebesgue measure of $\Omega$, while we let $|\Omega|_{*}=|\Omega\cap\mathbb{Z}^{d}|$ be the number of multi-integers contained in $\Omega$. It is not hard to show that
@@ 1.2 Some geometric concepts
We say $\mathcal{X}$ consists of clumps with parameters $(p,\tau,\lambda)$ if the following hold. There exist $r\in\mathbb{N}_{+}$ and disjoint subsets $\mathcal{C}_{1},\dots,\mathcal{C}_{r}$ of $\mathcal{X}$ called clumps such that $\mathcal{X}=\mathcal{C}_{1}\cup\cdots\cup\mathcal{C}_{r}$, the $\ell^{p}$ distance between any two elements in the same clump is bounded above by $\tau$, and $\max_{1\leq k\leq r}$ $|\mathcal{C}_{k}|=\lambda$. Further, if $r>1$, assume that the $\ell^{p}$ distance between different $\mathcal{C}_{j}$ and $\mathcal{C}_{k}$ exceeds $\tau$.
Aside from clumps, a more complicated geometric description can be given in terms of hyperplane decompositions. We say a set $\mathcal{X}$ consists of $r$ hyperplanes relative to a $x\in\mathcal{X}$ (called the reference point) if there exist hyperplanes $\mathcal{H}_{1},\dots,\mathcal{H}_{r}\subseteq\mathbb{R}^{d}$ that do not intersect $x$ and
@@ 1.3 Summary of main results and organization
There are many applications where a strong separation condition is violated. In Section 3, we focus on the super-resolution regime, which allows for $\Delta_{p^{\prime}}(\mathcal{X})$ to be arbitrarily small. As the separation goes to zero, the smallest singular values of both Fourier operators tend to zero, and our primary goal is to derive non-asymptotic and sharp lower bounds. Theorems 3.9 and 3.10 are multiscale estimates which show that the distances between elements in $\mathcal{X}$ at small scales predominately influence these operators’ smallest singular values. We specialize these theorems to clumps in order to give more interpretable results. Corollaries 3.11 and 3.12 show that if $\mathcal{X}$ consists of clumps with parameters $(p,\tau,\lambda)$ where $\tau\gtrsim\frac{\lambda}{m}$ and $\Delta_{p^{\prime}}(\mathcal{X})$ is sufficiently small, then
@@ 2 Well-separated case
Let $\Omega\subseteq\mathbb{R}^{d}$ be compact, $p\in[1,\infty]$, and $\delta\in(0,\frac{1}{2})$. Suppose there exist $\varphi$ and $\psi$ that are continuous on $\mathbb{R}^{d}$, supported in $\Omega_{\delta}^{p}$, $\widehat{\varphi},\widehat{\psi}\in L^{1}(\mathbb{R}^{d})$, and $\widehat{\psi}(\xi)\leq\mathbbm{1}_{\Omega}(\xi)\leq\widehat{\varphi}(\xi)$ for all $\xi\in\mathbb{R}^{d}$. For any set $\mathcal{X}\subseteq[-\frac{1}{2},\frac{1}{2})^{d}$ such that $\Delta_{p}(\mathcal{X})>\delta$, we have
@@
For any $\delta>0$ and integer $d\geq 2$, the following hold. There exist $\psi$ and $\varphi$ that are uniformly continuous on $\mathbb{R}^{d}$, supported in $Q_{\delta}$, $\widehat{\varphi},\widehat{\psi}\in L^{1}(\mathbb{R}^{d})$, $\widehat{\psi}(\xi)\leq\mathbbm{1}_{Q_{1}}(\xi)\leq\widehat{\varphi}(\xi)$ for all $\xi\in\mathbb{R}^{d}$, and
$\int_{\mathbb{R}^{d}}(\widehat{\varphi}-\mathbbm{1}_{Q_{1}})=\int_{\mathbb{R}^{d}}(\mathbbm{1}_{Q_{1}}-\widehat{\psi})$ $=\left(\big{(}1+\tfrac{1}{2\delta}\big{)}^{d}-1\right)\,|Q_{1}|.$
@@
$\psi(0)$ $=|Q_{1}|-\int_{\mathbb{R}^{d}}(\mathbbm{1}_{Q_{1}}-\widehat{\psi})=\left(2-\big{(}1+\tfrac{1}{2\delta}\big{)}^{d}\right)\,|Q_{1}|>(2-e^{d/(2\delta)})|Q_{1}|\geq 0,$ (2.1)
@@
Notice that the $\ell^{\infty}$ separation condition for $\mathcal{X}$ is inversely proportional to $m$, which is expected from scaling arguments. Additionally, the constant $\beta d$ grows linearly in the dimension $d$, but it is not clear whether this growth in $d$ is necessary.
Let us compare Theorem 2.3 with Kunis-Nagel-Strotmann [32, Theorem 4.2]. They parameterized the frequency set by $\{0,\dots,N-1\}^{d}$, while we use $Q_{m}=\{-m,\dots,m\}^{d}$. The singular values of $\Phi_{\Omega,\mathcal{X}}$ are invariant under multi-integer shifts of $\Omega$, so our settings are equivalent whenever we enforce the relationship $N=2m+1$. They proved that $\Delta_{2}(\mathcal{X})>\frac{4d}{2m+1}$ implies $\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})>0.9\,\sqrt{N^{d}}=0.9\sqrt{|Q_{m}|_{*}}$. To get the same conclusion, we pick $\beta=(2\log(2-0.81))^{-1}\approx 0.7671$ in Theorem 2.3 to see that $\Delta_{2}(\mathcal{X})\geq\frac{\beta d}{m}$ implies $\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})\geq 0.9\,\sqrt{|Q_{m}|_{*}}$. Both results require separation conditions that grow linearly in $d$, but our universal constant is smaller due to Barton.
@@
$c(\alpha):=\frac{1}{|\mathbb{S}^{d-1}|}\left(\frac{2}{\alpha}\right)^{d}\frac{\gamma(\alpha)}{1+\gamma(\alpha)/d}\quad\text{and}\quad\gamma(\alpha):=\frac{-\pi\alpha J_{d/2-1}(\pi\alpha)}{J_{d/2}(\pi\alpha)}.$ (2.3)
@@
Let $d\geq 2$ be an integer and $\alpha\in(\frac{j_{d/2-1,1}}{\pi},\frac{j_{d/2,1}}{\pi})$. Then there exists a uniformly continuous $\psi\colon\mathbb{R}^{d}\to\mathbb{R}$ supported in $B_{\alpha}$ such that $\widehat{\psi}\in L^{1}(\mathbb{R}^{d})$, $\widehat{\psi}(\xi)\leq\mathbbm{1}_{B_{1}}(\xi)$ for all $\xi\in\mathbb{R}^{d}$ and $\psi(0)=c(\alpha).$ If $\alpha\leq\frac{j_{d/2-1,1}}{\pi}$, there is no $\psi$ satisfying these properties with $\psi(0)>0$.
@@
$c(\alpha):=\frac{1}{|\mathbb{S}^{d-1}|}\left(\frac{2\pi}{j_{d/2,1}}\right)^{d}\quad\text{if}\quad\alpha\geq\frac{j_{d/2,1}}{\pi}.$ (2.4)
@@
There is a related but not directly comparable result. For $d=2$, Chen-Moitra [14] studied how small $\Delta_{2}(\mathcal{X})$ can be so that $\sigma_{\min}(\mathcal{F}_{B_{1/2},\mathcal{X}})$ is not exponentially small in $s$. By selecting $\mathcal{X}$ to be $s$ points on a hexagonal lattice of $\mathbb{R}^{2}$ and letting $s$ grow, they showed that $\Delta_{2}(\mathcal{X})\geq\sqrt{4/3}\approx 1.1547$ is required for stable recovery of $\mu$, in some appropriate sense defined in their paper. This construction can be adapted to derive an upper bound (as opposed to lower) for $\sigma_{\min}(\mathcal{F}_{B_{1/2},\mathcal{X}})$. Their conclusion that a minimum separation of $\sqrt{4/3}$ is necessary requires allowing $s$ to arbitrarily large. We will see in the next section that provided $s$ is fixed, the minimum separation can be arbitrarily small.
@@ 3 Super-resolution regime and arbitrary nodes
In the super-resolution regime, a primary method that we will use to lower bound $\sigma_{\min}(\mathcal{F}_{\Omega,\mathcal{X}})$ and $\sigma_{\min}(\Phi_{\Omega,\mathcal{X}})$ is through a “dual” relationship with minimum norm interpolation. This duality was introduced in [36, Proposition 2.12] for dealing with $\Phi_{\Omega,\mathcal{X}}$ in the one dimensional case. It readily generalizes to higher dimensions and to $\mathcal{F}_{\Omega,\mathcal{X}}$ with minor modifications. This dual relationship was further combined with density ideas and sparsity decompositions in [35], which significantly improved existing estimates for one-dimensional Fourier matrices.
We first collect these ideas before moving onto the main theorems. While we provide some exposition, it is rather brief; for those interested in the intuition and concepts behind these techniques, we refer the reader to [35], which explains them in greater detail and in the simpler one-dimensional case. Recall that throughout, we assume $\Omega\subseteq\mathbb{R}^{d}$ is compact. We let $\mathcal{W}(\Omega)$ denote the space of bandlimited functions or the Paley-Wiener space; that is, a complex-valued $f$ belongs to $\mathcal{W}(\Omega)$ if and only if there is a $F\in L^{2}(\Omega)$ such that
@@
This is a straightforward adaptation of [36, Proposition 2.12], but we include its proof in Section 6.2 for completeness. Typically, we will use this proposition for the case when $\varepsilon=0$ and $v$ is a right singular vector of $\mathcal{F}_{\Omega,\mathcal{X}}$ (or $\Phi_{\Omega,\mathcal{X}}$) corresponding to its smallest singular value. It is prudent to mention that this proposition does not make any claims about the relationship between singular values $\mathcal{F}_{\Omega,\mathcal{X}}$ and $\Phi_{\Omega,\mathcal{X}}$ because they in general have different right singular vectors, so this statement would be applied to different $v$’s.
@@
The next proposition acts as a converse to Proposition 3.1, and shows that any lower bound on the smallest singular value provides the existence of polynomials with prescribed interpolation properties. A proof of the following lemma for $d=1$ can be found in [35, Proposition 5.3] and readily extends to higher dimensions.
@@
This proposition can be readily deduced from a greedy construction of the $\mathcal{X}_{1},\dots,\mathcal{X}_{\nu}$, see the proof of [35, Proposition 6.1] for $d=1$. The next proposition illustrates a nontrivial application of the polynomial method and is proved in Section 6.4.
@@
$\sigma_{\min}\left(\Phi_{\Omega^{p}_{C/\tau},\,\mathcal{W}}\right)\geq c\sqrt{|\Omega^{p}_{C/\tau}|_{*}}.$ (3.2)
@@
Provided that the assumptions hold, this localization comes at a price of course. Comparing both sides of the inequality in (3.3), the number of Fourier samples are reduced from $|\Omega^{p}_{m}|_{*}$ to $|\Omega^{p}_{m/2}|_{*}$, so resolution worsens by a factor of 2. We also pick up an additional $c^{\nu_{p}(\tau,\mathcal{X})}$ term, which can be small if either $c$ is close to zero and/or $\nu_{p}(\tau,\mathcal{X})$ is large. For many sets of interest, such as clumps, the $\tau$ local sparsity of $\mathcal{X}$ is small. When we apply this theorem, the constant $c$ will come from Theorems 2.5 and 2.3, so it is important to obtain good constants for the well separated case to temper the effects of localization.
With the localization lemma at hand, let us explain at a high level how to deal with the local set $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$. We will again employ the duality principle, which transforms the problem to constructing Lagrange polynomials $\{f_{k}\}_{k=1}^{s}$ for $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$. Shifting by $x_{k}$ and noting that both $\mathcal{P}(\Omega)$ and $\mathcal{W}(\Omega)$ are invariant under shifts, this reduces the problem down to the canonical case whereby $0\in\mathcal{U}$ and all other elements in $\mathcal{U}$ are sufficiently close to zero. Importantly, we place no lower bounds on $\Delta_{p}(\mathcal{U})$. We seek a $f$ such that $f(0)=1$, $f$ vanishes on $\mathcal{U}\setminus\{0\}$, and its $L^{2}$ norm is suitably controlled.
@@
Second, we need to pick $q(u)$ carefully so that $|q(u)\cdot u|$ is not too close to zero, in order to ensure that $\|\widetilde{f}\|_{L^{\infty}(\mathbb{T}^{d})}$ is reasonably controlled, yet $|q(u)|_{p}$ itself is not too large otherwise the obtained interpolant might not be in $\mathcal{P}(\Omega)$. See Figure 2 for a cartoon of how replacing $a$ with $q$ effectively changes the direction of $f$.
@@
$\sigma_{\min}(\mathcal{F}_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{2^{\nu-1}s}}\left(\frac{c(\alpha)}{|B_{1}|}\right)^{\nu/2}\left(\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\right)^{\nu/2}\,\sqrt{|B_{m/(2\nu)}|}\,\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{2}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{2}\Bigg{\}},$
$\sigma_{\min}(\Phi_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{2^{\nu-1}s}}\left(\frac{c(\alpha)}{|B_{1}|}\right)^{\nu/2}\left(\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\right)^{\nu/2}\sqrt{|B_{m/(2\nu)}|_{*}}\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{2}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{2}\Bigg{\}}.$
@@
$\sigma_{\min}(\mathcal{F}_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\left(1-\frac{1}{2}e^{1/(2\beta)}\right)^{\nu/2}\,\sqrt{|Q_{m/(2\nu)}|}\,\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{1}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{1}\Bigg{\}},$
$\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\left(1-\frac{1}{2}e^{1/(2\beta)}\right)^{\nu/2}\sqrt{|Q_{m/(2\nu)}|_{*}}\,\min_{1\leq k\leq s}\Bigg{\{}\prod_{0<|x_{j}-x_{k}|_{1}\leq\frac{\nu}{m}}\,\frac{m}{\nu}\,|x_{j}-x_{k}|_{1}\Bigg{\}}.$